OpenStudy (anonymous):

(a+b)^3/(x-y)^2 * (x-y)/(5-p) * (p-5)^2/(a+b)^2=

OpenStudy (anonymous):

\[ \dfrac{(a+b)^3}{(x-y)^2}\cdot\dfrac{(x-y)}{(5-p)}\cdot\dfrac{(p-5)^2}{(a+b)^2}\\ \dfrac{(a+b)}{(x-y)}\cdot\dfrac{1}{-(p-5)}\cdot\dfrac{(p-5)^2}{1}\\ \dfrac{-(a+b)(p-5)}{(x-y)}\\ \]

OpenStudy (anonymous):

I see. thank you very much!

OpenStudy (anonymous):

actually where did the negative on top come from? the answer is (a+b)(5-p)/(x-y)